# Frobenius pull-back of parabolic bundles and dormant opers

### v1

@@ Abstract.
footnotetext: Y. Wakabayashi: Graduate School of Information Science and Technology, Osaka University, Suita, Osaka 565-0871, Japan;footnotetext: e-mail: wakabayashi@ist.osaka-u.ac.jp;footnotetext: 2020 Mathematical Subject Classification: Primary 14H60, Secondary 14G17;footnotetext: Key words: parabolic bundle, positive characteristic, Frobenius pull-back, dormant oper, Frobenius-destabilized bundle
@@ 1. Introduction
Let $(g,r)$ be a pair of nonnegative integers with $02g-2+r>$ 0 and $\mathscr{X}:=(X,\{\sigma_{i}\}_{i=1}^{r})$ an $r$-pointed smooth proper curve of genus $g$ over an algebraically closed field $k$. A parabolic (vector) bundle is a vector bundle equipped with weighted flags on its fibers over the marked points $\sigma_{i}$. To be more precise, a parabolic bundle on $\mathscr{X}$ means a collection
$\mathscr{F}:=(\mathcal{F},\vec{\mathpzc{f}},\vec{\alpha})$ (1.1)
(cf. Definition 2.3) consisting of a vector bundle (i.e., a locally free coherent sheaf) $\mathcal{F}$ on $X$, a set of parabolic weights $\vec{\alpha}$, and an $r$-tuple $(\mathpzc{f}_{1},\cdots,\mathpzc{f}_{r})$, where $\mathpzc{f}_{i}$ ($i=1,\cdots,r$) denotes a sequence of $k$-linear surjections
@@
$\left(\begin{matrix}\text{the category of }\\ \text{vector bundles on $X^{(1)}$}\end{matrix}\right)\xrightarrow{\sim}\left(\begin{matrix}\text{the category of}\\ \text{$p$-flat bundles on $X$}\end{matrix}\right),$ (1.3)
@@
where $a_{i}^{[j]}$’s and $\ell_{i}^{[j]}$’s are the parabolic weights and types, respectively, of the parabolic structure $(\vec{\mathpzc{f}},\vec{\alpha})$ at $\sigma_{i}$ (cf. (2.9)). Then, we say that $\mathscr{F}$ is semistable (resp., stable) if the inequality $\mathrm{par}\text{-}\mu(\mathscr{E})\leq\mathrm{par}\text{-}\mu(\mathscr{F})$ (resp., $\mathrm{par}\text{-}\mu(\mathscr{E})<\mathrm{par}\text{-}\mu(\mathscr{F})$) holds for any nontrivial proper parabolic subbundle $\mathscr{E}$ of $\mathscr{F}$ (cf. Definition 2.6).
@@
To state our result concerning its parabolic (and higher-level) generalization, we prepare some notations. Let $\mathcal{L}$ be a line bundle on $X^{(N)}$ and $\vec{a}$ an element of $(\Xi_{n,N}^{<})^{\times r}$ such that $s_{1}^{1}(\vec{a})\in(\Xi_{n,1}^{\leq})^{\times r}$ and $s_{2}^{1}(\vec{a})\in(\Xi_{n,N-1}^{\leq})^{\times r}$ (cf. (4.25) for the definitions of $s_{1}^{1}(-)$ and $s_{2}^{1}(-)$). We set $\vec{1}_{\times n}:=(1_{\times n},\cdots,1_{\times n})$ (= the $r$-tuple of copies of $1_{\times n}:=(1,\cdots,1)\in\Xi_{n,N}^{\leq}$), $\omega:=(n,\vec{1}_{\times n},\vec{a}/p^{N})$, and $\mathcal{L}^{F}_{\flat}:=((F^{(N)*}_{X/k}\mathcal{L})(D_{\star}),\nabla^{\mathrm{can}}_{\mathcal{L},D_{\star}})$ (cf. (3.3)), where $D_{\star}$ denotes the divisor $\sum\limits_{i=1}^{r}(\sum\limits_{j=1}^{n}a_{i}^{[j]})\sigma_{i}$. Then, we obtain the moduli category
$\mathcal{U}^{F}_{\mathscr{X}^{(N)},n,\omega,\mathcal{L}}\ \left(\text{resp.,}\ \mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}^{F}}\right)$ (1.7)
@@
  • $\sum\limits_{i=1}^{r}(s_{1}^{1}(a_{i}^{[n]})-s_{1}^{1}(a_{i}^{[1]}))<\frac{n(2g-2+r)}{2}\leq\frac{p}{n}$.
@@
Let us take a line bundle $\mathcal{L}$ on $X^{(N)}$ of even degree and an element $\vec{a}:=(a_{1},\cdots,a_{r})\in(\Xi_{2,1}^{<})^{\times r}$, where $a_{i}:=(a_{i}^{[1]},a_{i}^{[2]})$. We write $\omega:=(2,\vec{1}_{\times 2},\vec{a}/p)$. Suppose further that the following two conditions are fulfilled:
  • $r+\sum\limits_{i=1}^{r}(a_{i}^{[1]}+a_{i}^{[2]})$ is even and $\sum\limits_{i=1}^{r}(a_{i}^{[2]}-a_{i}^{[1]})<2g-2+r\leq\frac{p}{2}$;
@@
Since $X^{\mathrm{log}}/S$ is log smooth, the sheaf of logarithmic $1$-forms $\Omega_{X^{\mathrm{log}}/S}$ of $X^{\mathrm{log}}/S$, as well as its dual $\mathcal{T}_{X^{\mathrm{log}}/S}:=\Omega_{X^{\mathrm{log}}/S}^{\vee}$, is a line bundle. In fact, we have natural isomorphisms $\Omega_{X^{\mathrm{log}}/S}\cong\Omega_{X/S}(D)$, $\mathcal{T}_{X^{\mathrm{log}}/S}\cong\mathcal{T}_{X/S}(-D)$. When there is no fear of confusion, we write $\Omega$ and $\mathcal{T}$ instead of $\Omega_{X^{\mathrm{log}}/S}$ and $\mathcal{T}_{X^{\mathrm{log}}/S}$, respectively. Also, for a positive integer $N$, we denote by $\mathcal{D}^{(N-1)}_{X^{\mathrm{log}}/S}$ (cf.  [Mon, Definition 2.3.1]) the sheaf of logarithmic differential operators on $X^{\mathrm{log}}/S$ of level $N-1$.
@@ 2.1. Definition of a parabolic bundle
    $\mathpzc{f}_{i}:\mathcal{F}_{i}=\mathcal{F}_{i}^{[m_{i}]}\xrightarrow{\mathpzc{f}_{i}^{[m_{i}]}}\mathcal{F}_{i}^{[m_{i}-1]}\xrightarrow{\mathpzc{f}_{i}^{[m_{i}-1]}}\cdots\xrightarrow{\mathpzc{f}_{i}^{[2]}}\mathcal{F}_{i}^{[1]}\xrightarrow{\mathpzc{f}_{i}^{[1]}}\mathcal{F}_{i}^{[0]}=0$ (2.1)
@@
    $(\mathpzc{f}_{i},\alpha_{i})$ (2.2)
    consisting of a quasi-parabolic structure $\mathpzc{f}_{i}$ as in (2.1) and an $m_{i}$-tuple of real numbers $\alpha_{i}:=(\alpha^{[1]}_{i},\cdots,\alpha^{[m_{i}]}_{i})$ with $0\leq\alpha^{[1]}_{i}\leq\cdots\leq\alpha^{[m_{i}]}_{i}$.
@@
    $\vec{\mathpzc{f}}\ \left(\text{resp.,}\ (\vec{\mathpzc{f}},\vec{\alpha})\right)$ (2.3)
    consisting of an $r$-tuple $\vec{\mathpzc{f}}:=(\mathpzc{f}_{1},\cdots,\mathpzc{f}_{r})$ (resp., $r$-tuples $\vec{\mathpzc{f}}:=(\mathpzc{f}_{1},\cdots,\mathpzc{f}_{r})$ and $\vec{\alpha}:=(\alpha_{1},\cdots,\alpha_{r})$) such that, for each $i=1,\cdots,r$, $\mathpzc{f}_{i}$ (resp., $(\mathpzc{f}_{i},\alpha_{i})$) forms a quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$ at $\sigma_{i}$.
@@
    Let $i$ be an element of $\{1,\cdots,r\}$ and $\mathpzc{f}_{i}$ (resp., $(\mathpzc{f}_{i},\alpha_{i})$) a quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$ at $\sigma_{i}$ as in (2.2). Then, the integer $m_{i}$ is called the length of $\mathpzc{f}_{i}$ (resp., $(\mathpzc{f}_{i},\alpha_{i})$). Also, if $S$ is connected, then we obtain the $m_{i}$-tuple of nonnegative integers
@@
    where $\ell_{i}^{[j]}$ denotes the rank of the vector bundle $\mathrm{Ker}(\mathpzc{f}_{i}^{[j]})$, and it is called the type of $\mathpzc{f}_{i}$ (resp., $(\mathpzc{f}_{i},\alpha_{i})$).
@@
    Let $\vec{\mathpzc{f}}$ (resp., $(\vec{\mathpzc{f}},\vec{\alpha})$), where $\vec{\mathpzc{f}}:=(\mathpzc{f}_{1},\cdots,\mathpzc{f}_{r})$ (resp., $\vec{\mathpzc{f}}:=(\mathpzc{f}_{1},\cdots,\mathpzc{f}_{r})$ and $\vec{\alpha}:=(\alpha_{1},\cdots,\alpha_{r})$), be a quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$. Under the assumption that $S$ is connected, we denote by $\ell_{i}$ the type of $\mathpzc{f}_{i}$ (resp., $(\mathpzc{f}_{i},\alpha_{i})$). Then, the collections
@@
    are called the length and type of $\vec{\mathpzc{f}}$ (resp., $(\vec{\mathpzc{f}},\vec{\alpha})$), respectively.
@@
    $\mathscr{F}:=(\mathcal{F},\vec{\mathpzc{f}})\ \left(\text{resp.,}\ \mathscr{F}:=(\mathcal{F},\vec{\mathpzc{f}},\vec{\alpha})\right),$ (2.6)
    consisting of a vector bundle $\mathcal{F}$ on $X$ of constant rank and a quasi-parabolic (resp., parabolic) structure $\vec{\mathpzc{f}}$ (resp., $(\vec{\mathpzc{f}},\vec{\alpha})$) of $\mathcal{F}$. If $\mathpzc{f}_{i}:=\{\mathpzc{f}_{i}^{[j^{\prime}]}\}_{j^{\prime}=1}^{m_{i}}$ denotes the $i$-th factor of $\vec{\mathpzc{f}}$ (where $i\in\{1,\cdots,r\}$), then, for each $j=1,\cdots,m_{i}$, we obtain a composite surjection
    $\varpi_{\mathscr{F},i}^{[j]}:\mathcal{F}\twoheadrightarrow\sigma_{i*}\mathcal{F}_{i}^{[m_{i}]}\xrightarrow{\sigma_{i*}\mathpzc{f}_{i}^{[m_{i}]}}\cdots\xrightarrow{\sigma_{i*}\mathpzc{f}_{i}^{[j+1]}}\sigma_{i*}\mathcal{F}_{i}^{[j]}.$ (2.7)
@@
Next, let $\mathscr{F}:=(\mathcal{F},\vec{\mathpzc{f}},\vec{\alpha})$ and $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{\alpha})$ be two parabolic bundles on $\mathscr{X}$ with parabolic structures having common length and parabolic weights
@@ 2.2. Stability for parabolic bundles
Let $\mathscr{F}:=(\mathcal{F},\vec{\mathpzc{f}},\vec{\alpha})$ be a parabolic bundle on $\mathscr{X}$ such that $(\vec{\mathpzc{f}},\vec{\alpha})$ is a collection as in the resp’d portion of (2.3). Given a subbundle $\mathcal{E}$ of $\mathcal{F}$ of constant rank, we set $\mathcal{F}^{[j]}_{i,\mathcal{E}}$ (for each $i=1,\cdots,r$ and $j=1,\cdots,m_{i}$) to be the image of $\mathcal{E}$ via $\varpi_{\mathscr{F},i}^{[j]}$. Assume that $\mathcal{F}^{[j]}_{i,\mathcal{E}}$ are all vector bundles on $S$ when restricted via $\sigma_{i}$. (Of course, this assumption is automatically satisfied if $S=\mathrm{Spec}(k)$.) Since $\mathpzc{f}_{i}^{[j]}$ restricts to a surjection $\mathpzc{f}_{i,\mathcal{E}}^{[j]}:\mathcal{F}^{[j]}_{i,\mathcal{E}}\twoheadrightarrow\mathcal{F}^{[j-1]}_{i,\mathcal{E}}$, the set of morphisms $\mathpzc{f}_{i,\mathcal{E}}:=\{\mathpzc{f}_{i,\mathcal{E}}^{[j]}\}_{j=1}^{m_{i}}$ specifies, via $\sigma_{i}^{*}(-)$, a quasi-parabolic structure on $\mathcal{E}$ at $\sigma_{i}$. The resulting collection $\mathscr{E}:=(\mathcal{E},\vec{\mathpzc{f}}_{\mathcal{E}},\vec{\alpha})$, where $\vec{\mathpzc{f}}_{\mathcal{E}}:=(\mathpzc{f}_{1,\mathcal{E}},\cdots,\mathpzc{f}_{r,\mathcal{E}})$, forms a parabolic bundle on $\mathscr{X}$; such a parabolic bundle obtained in this way is called parabolic subbundle of $\mathscr{F}$.
@@
where $\mathrm{deg}(\mathcal{E})$ denotes the relative degree of $\mathcal{E}$ and $\ell^{[j]}_{i,\mathcal{E}}$ (for each $i$ and $j$) denotes the rank of $\mathrm{Ker}(\mathpzc{f}^{[j]}_{i,\mathcal{E}})$ (viewed as a vector bundle on $S$). We refer to $\mathrm{par}\text{-}\mathrm{deg}(\mathscr{E})$ and $\mathrm{par}\text{-}\mu(\mathscr{E})$ as, respectively, the parabolic degree and the parabolic slope of $\mathscr{E}$. The formations of $\mathrm{par}\text{-}\mathrm{deg}(-)$ and $\mathrm{par}\text{-}\mu(-)$ commute with the base-change over any $S$-scheme.

### v2

@@ Abstract.
footnotetext: Y. Wakabayashi: Graduate School of Information Science and Technology, The University of Osaka, Suita, Osaka 565-0871, Japan;footnotetext: e-mail: wakabayashi@ist.osaka-u.ac.jp;footnotetext: 2020 Mathematical Subject Classification: Primary 14H60, Secondary 14G17;footnotetext: Key words: parabolic bundle, positive characteristic, Frobenius pull-back, dormant oper, Frobenius-destabilized bundle
@@ 1. Introduction
Let $(g,r)$ be a pair of nonnegative integers with 2g2+r>02g-2+r>0 and $\mathscr{X}:=(X,\{\sigma_{i}\}_{i=1}^{r})$ an $r$-pointed smooth proper curve of genus $g$ over an algebraically closed field $k$. A parabolic (vector) bundle is a vector bundle equipped with weighted flags on its fibers over the marked points $\sigma_{i}$. To be more precise, a parabolic bundle on $\mathscr{X}$ means a collection
$\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha})$ (1.1)
(cf. Definition 2.3) consisting of a vector bundle (i.e., a locally free coherent sheaf) $\mathcal{F}$ on $X$, a set of parabolic weights $\vec{\alpha}$, and an $r$-tuple $(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$, where $\mathfrak{f}_{i}$ ($i=1,\cdots,r$) denotes a sequence of $k$-linear surjections
@@
[…]
@@
where $a_{i}^{[j]}$’s and $\ell_{i}^{[j]}$’s are the parabolic weights and types, respectively, of the parabolic structure $(\vec{\mathfrak{f}},\vec{\alpha})$ at $\sigma_{i}$ (cf. (2.9)). Then, we say that $\mathscr{F}$ is semistable (resp., stable) if the inequality $\mathrm{par}\text{-}\mu(\mathscr{E})\leq\mathrm{par}\text{-}\mu(\mathscr{F})$ (resp., $\mathrm{par}\text{-}\mu(\mathscr{E})<\mathrm{par}\text{-}\mu(\mathscr{F})$) holds for any nontrivial proper parabolic subbundle $\mathscr{E}$ of $\mathscr{F}$ (cf. Definition 2.6).
@@
To state our result concerning its parabolic (and higher-level) generalization, we prepare some notations. Let $n$ be a positive integer, $\mathcal{L}$ a line bundle on $X^{(N)}$, and $\vec{a}$ an element of $(\Xi_{n,N}^{<})^{\times r}$ as in § 5.1 such that $s_{1}^{1}(\vec{a})\in(\Xi_{n,1}^{\leq})^{\times r}$ and $s_{2}^{1}(\vec{a})\in(\Xi_{n,N-1}^{\leq})^{\times r}$ (cf. (4.25) for the definitions of $s_{1}^{1}(-)$ and $s_{2}^{1}(-)$). We set $\vec{1}_{\times n}:=(1_{\times n},\cdots,1_{\times n})$ (= the $r$-tuple of copies of $1_{\times n}:=(1,\cdots,1)\in\Xi_{n,N}^{\leq}$), $\omega:=(n,\vec{1}_{\times n},\vec{a}/p^{N})$, and $\mathcal{L}^{F}_{\flat}:=((F^{(N)*}_{X/k}\mathcal{L})(D_{\star}),\nabla^{\mathrm{can}}_{\mathcal{L},D_{\star}})$ (cf. (3.3)), where $D_{\star}$ denotes the divisor $\sum\limits_{i=1}^{r}(\sum\limits_{j=1}^{n}a_{i}^{[j]})\sigma_{i}$. Then, we obtain the moduli category
$\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}\ \left(\text{resp.,}\ \mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}^{F}}\right)$ (1.7)
@@
  • $\sum\limits_{i=1}^{r}(s_{1}^{1}(a_{i}^{[n]})-s_{1}^{1}(a_{i}^{[1]}))<\frac{n(2g-2+r)}{2}<\frac{p}{n}$.
@@
Let us take a line bundle $\mathcal{L}$ on $X^{(1)}$ of even degree and an element $\vec{a}:=(a_{1},\cdots,a_{r})\in(\Xi_{2,1}^{<})^{\times r}$, where $a_{i}:=(a_{i}^{[1]},a_{i}^{[2]})$. We write $\omega:=(2,\vec{1}_{\times 2},\vec{a}/p)$. Suppose further that the following two conditions are fulfilled:
  • $r+\sum\limits_{i=1}^{r}(a_{i}^{[1]}+a_{i}^{[2]})$ is even and $\sum\limits_{i=1}^{r}(a_{i}^{[2]}-a_{i}^{[1]})<2g-2+r<\frac{p}{2}$;
@@
Since $X^{\mathrm{log}}/S$ is log smooth, the sheaf of logarithmic $1$-forms $\Omega_{X^{\mathrm{log}}/S}$ of $X^{\mathrm{log}}/S$, as well as its dual $\mathcal{T}_{X^{\mathrm{log}}/S}:=\Omega_{X^{\mathrm{log}}/S}^{\vee}$, is a line bundle. In fact, we have natural isomorphisms $\Omega_{X^{\mathrm{log}}/S}\cong\Omega_{X/S}(D)$, $\mathcal{T}_{X^{\mathrm{log}}/S}\cong\mathcal{T}_{X/S}(-D)$. When there is no fear of confusion, we write $\Omega$ and $\mathcal{T}$ instead of $\Omega_{X^{\mathrm{log}}/S}$ and $\mathcal{T}_{X^{\mathrm{log}}/S}$, respectively. Also, for a positive integer $N$, we denote by $\mathcal{D}^{(N-1)}_{X^{\mathrm{log}}/S}$ (cf.  [Mon, Definition 2.3.1]) the sheaf of logarithmic differential operators on $X^{\mathrm{log}}/S$ of level N1N-1.
@@ 2.1. Definition of a parabolic bundle
    $\mathfrak{f}_{i}:\mathcal{F}_{i}=\mathcal{F}_{i}^{[m_{i}]}\xrightarrow{\mathfrak{f}_{i}^{[m_{i}]}}\mathcal{F}_{i}^{[m_{i}-1]}\xrightarrow{\mathfrak{f}_{i}^{[m_{i}-1]}}\cdots\xrightarrow{\mathfrak{f}_{i}^{[2]}}\mathcal{F}_{i}^{[1]}\xrightarrow{\mathfrak{f}_{i}^{[1]}}\mathcal{F}_{i}^{[0]}=0$ (2.1)
@@
    $(\mathfrak{f}_{i},\alpha_{i})$ (2.2)
[…]
@@
    $\vec{\mathfrak{f}}\ \left(\text{resp.,}\ (\vec{\mathfrak{f}},\vec{\alpha})\right)$ (2.3)
    consisting of an $r$-tuple $\vec{\mathfrak{f}}:=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$ (resp., $r$-tuples $\vec{\mathfrak{f}}:=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$ and $\vec{\alpha}:=(\alpha_{1},\cdots,\alpha_{r})$) such that, for each $i=1,\cdots,r$, $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$) forms a quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$ at $\sigma_{i}$.
@@
    Let $i$ be an element of $\{1,\cdots,r\}$ and $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$) a quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$ at $\sigma_{i}$ as in (2.2). Then, the integer $m_{i}$ is called the length of $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$). Also, if $S$ is connected, then we obtain the $m_{i}$-tuple of nonnegative integers
@@
    where $\ell_{i}^{[j]}$ denotes the rank of the vector bundle $\mathrm{Ker}(\mathfrak{f}_{i}^{[j]})$, and it is called the type of $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$).
@@
    Let $\vec{\mathfrak{f}}$ (resp., $(\vec{\mathfrak{f}},\vec{\alpha})$), where $\vec{\mathfrak{f}}:=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$ (resp., $\vec{\mathfrak{f}}:=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$ and $\vec{\alpha}:=(\alpha_{1},\cdots,\alpha_{r})$), be a quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$. Under the assumption that $S$ is connected, we denote by $\ell_{i}$ the type of $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$). Then, the collections
@@
    are called the length and type of $\vec{\mathfrak{f}}$ (resp., $(\vec{\mathfrak{f}},\vec{\alpha})$), respectively.
@@
    $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}})\ \left(\text{resp.,}\ \mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha})\right),$ (2.6)
    consisting of a vector bundle $\mathcal{F}$ on $X$ of constant rank and a quasi-parabolic (resp., parabolic) structure $\vec{\mathfrak{f}}$ (resp., $(\vec{\mathfrak{f}},\vec{\alpha})$) of $\mathcal{F}$. If $\mathfrak{f}_{i}:=\{\mathfrak{f}_{i}^{[j^{\prime}]}\}_{j^{\prime}=1}^{m_{i}}$ denotes the $i$-th factor of $\vec{\mathfrak{f}}$ (where $i\in\{1,\cdots,r\}$), then, for each $j=1,\cdots,m_{i}$, we obtain a composite surjection
    $\varpi_{\mathscr{F},i}^{[j]}:\mathcal{F}\twoheadrightarrow\sigma_{i*}\mathcal{F}_{i}^{[m_{i}]}\xrightarrow{\sigma_{i*}\mathfrak{f}_{i}^{[m_{i}]}}\cdots\xrightarrow{\sigma_{i*}\mathfrak{f}_{i}^{[j+1]}}\sigma_{i*}\mathcal{F}_{i}^{[j]}.$ (2.7)
@@
Next, let $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha})$ and $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{\alpha})$ be two parabolic bundles on $\mathscr{X}$ with parabolic structures having common length and parabolic weights
@@ 2.2. Stability for parabolic bundles
Let $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha})$ be a parabolic bundle on $\mathscr{X}$ such that $(\vec{\mathfrak{f}},\vec{\alpha})$ is a collection as in the resp’d portion of (2.3). Given a subbundle $\mathcal{E}$ of $\mathcal{F}$ of constant rank, we set $\mathcal{F}^{[j]}_{i,\mathcal{E}}$ (for each $i=1,\cdots,r$ and $j=1,\cdots,m_{i}$) to be the image of $\mathcal{E}$ via $\varpi_{\mathscr{F},i}^{[j]}$. Assume that $\mathcal{F}^{[j]}_{i,\mathcal{E}}$ are all vector bundles on $S$ when restricted via $\sigma_{i}$. (Of course, this assumption is automatically satisfied if $S=\mathrm{Spec}(k)$.) Since $\mathfrak{f}_{i}^{[j]}$ restricts to a surjection $\mathfrak{f}_{i,\mathcal{E}}^{[j]}:\mathcal{F}^{[j]}_{i,\mathcal{E}}\twoheadrightarrow\mathcal{F}^{[j-1]}_{i,\mathcal{E}}$, the set of morphisms $\mathfrak{f}_{i,\mathcal{E}}:=\{\mathfrak{f}_{i,\mathcal{E}}^{[j]}\}_{j=1}^{m_{i}}$ specifies, via $\sigma_{i}^{*}(-)$, a quasi-parabolic structure on $\mathcal{E}$ at $\sigma_{i}$. The resulting collection $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{f}}_{\mathcal{E}},\vec{\alpha})$, where $\vec{\mathfrak{f}}_{\mathcal{E}}:=(\mathfrak{f}_{1,\mathcal{E}},\cdots,\mathfrak{f}_{r,\mathcal{E}})$, forms a parabolic bundle on $\mathscr{X}$; such a parabolic bundle obtained in this way is called parabolic subbundle of $\mathscr{F}$.
@@
[…]